How many kg of corn that cost P40 per kg must be mixed with 10kg of peanut that cost P80 per kg to make a mixture that cost P55 per kg?
Answer by ptaylor(2198)

Let x=amount of corn needed
Now we know that the value of the corn (40x) plus the value fo the peanuts (10*80) before they are mixed has to equal the value of the final mixture (55*(10+x)). So our equation to solve is:
40x+10*80=55(10+x) get rid of parens
40x+800=550+55x subtract 800 and also 55x from each side
40x-55x+800-800=550-800+55x-55x collect like terms
-15x=-250 divide each side by -15
x=16.667 kg----------------------------amount of corn needed
CK
16.667*40+10*80=55*26.667
666.68+800=1466.685
1466.68=1466.68

Does this help???-----ptaylor